# Fully Dynamic Spectral Sparsification for Directed Hypergraphs

### v1

@@ 1.1 Related Work
We briefly discuss the related work on graphs and hypergraphs for several computational models.
@@
Starting with [ST11], spectral sparsification has been extensively studed on graphs [BSS12, KP12, BSST13, ZLO15, KLP16, LS17, LS18]. Recently, the concept has been extended to undirected and directed hypergraphs [SY19, BST19, KKTY21, KKTY21a, RY22, OST23, Lee23, JLS23, KPS24].
@@
For undirected graphs, there have been several results for dynamic cut sparsifiers that use a similar approach to ours. This includes [ADKKP16], which achieves polylogarithmic update time using $t$-bundle spanners, and its extension against an adaptive adversary by [BBGNSSS22] and to directed graphs [Zha25]. More recently, [KLP25] achieved a fully dynamic spectral sparsifier with near-optimal size and update time for undirected hypergraphs. More broadly, spectral-based sparsification is closely related to dynamic vertex sparsifiers [GHP18, DGGP19, CGHPS20, GLP21, AMV21, BGJLLPS22, DGGLPSY22].
@@
[KX16] achieved simple yet powerful algorithms for spectral graph sparsification in both parallel and distributed settings. Very recently, [GK25] developed a parallel batch-dynamic algorithm for spanners. Other related works include distributed vertex sparsifiers [ZLB21, FGLPSY21].
@@
For graphs, [KL13] extended the sampling scheme based on effective resistances [SS08] to the semi-streaming setting. Moreover, [CMP20] obtained an online spectral sparsification algorithm for graphs. Recently, [STY24] proposed an online algorithm for spectral hypergraph sparsification. For graphs in dynamic streams, [AGM13] developed a spectral sparsifier. Also, there has been a series of work on spectral sparsification in dynamic streams for both graphs and hypergraphs [KLMMS17, KNST19, KMMMNST20].
@@ 2 Order the hyperedges of $H$ by their weights in decreasing order
6
7foreach hyperedge $e$ of $H\setminus C$ do /* compute $S$ */
@@ 26   Pass the deletion of $e$ to $\mathcal{A}_{j}$
27
28
29Procedure OutputSparsifier()
30 return $\widetilde{H}=\bigcup_{i=1}^{i_{\text{last}}}\widetilde{H}_{i}$

### v2

@@ 1.1 Related Work
We briefly discuss the related work for spectral sparsification on both graphs and hypergraphs below.
@@
Starting with [ST11] [ST11], spectral sparsification has been extensively studed on graphs [BSS12, KP12, BSST13, ZLO15, KLP16, LS17, LS18]. Recently, the concept has been extended to undirected and directed hypergraphs [SY19, BST19, KKTY21, KKTY21a, RY22, OST23, Lee23, JLS23, KPS24].
@@
For undirected graphs, there have been several results for dynamic spectral sparsifiers that use a similar approach to ours. This includes the work of [ADKKP16] [ADKKP16], which achieves polylogarithmic update time using $t$-bundle spanners, and its extension against an adaptive adversary by [BBGNSSS22] [BBGNSSS22] and to directed graphs by [Zha25] [Zha25]. More recently, [KLP25] [KLP25] achieved a fully dynamic spectral sparsifier with near-optimal size and update time for undirected hypergraphs. A closely related notion of sparsification, namely vertex sparsifiers, has also been studied in the dynamic setting [GHP18, DGGP19, CGHPS20, GLP21, AMV21, BGJLLPS22, DGGLPSY22].
@@
[KX16] [KX16] achieved simple algorithms for spectral graph sparsification that can be implemented in many computational models, including both parallel and distributed settings, with sub-optimal guarantees on the sparsifier size. Very recently, [GK25] [GK25] developed a parallel batch-dynamic algorithm for spanners. Other related works include distributed vertex sparsifiers [ZLB21, FGLPSY21].
@@
For graphs, [KL13] [KL13] extended the sampling scheme based on effective resistances [SS08] to the semi-streaming setting. [CMP20] [CMP20] obtained an online spectral sparsification algorithm for graphs. Recently, [STY24] [STY24] proposed an online algorithm for spectral hypergraph sparsification. For graphs in dynamic streams, [AGM13] [AGM13] developed a spectral sparsifier. Also, there has been a series of work on spectral sparsification in dynamic streams for both graphs and hypergraphs [KLMMS17, KNST19, KMMMNST20].
@@ 2 Order the hyperedges of $H$ by their weights in decreasing order
6 end foreach
7
8foreach hyperedge $e$ of $H\setminus C$ do /* compute $S$ */
@@ 30   Pass the deletion of $e$ to $\mathcal{A}_{j}$
33
34Procedure OutputSparsifier()
35 return $\widetilde{H}=\bigcup_{i=1}^{i_{\text{last}}}\widetilde{H}_{i}$
36
37
38